# cp3-bench: A tool for benchmarking symbolic regression algorithms tested with cosmology

### v1

@@ Abstract
We present a benchmark study of ten symbolic regression algorithms applied to cosmological datasets. We find that the dimension of the feature space as well as precision of datasets are highly important for symbolic regression tasks to be successful. We find no indication that inter-dependence of features in datasets are particularly important, meaning that it is not an issue if datasets e.g. contain both $z$ and $H(z)$ as features. We find no indication that performance of algorithms on standardized datasets are good indications of performance on cosmological datasets. This suggests that it is not necessarily prudent to choose which symbolic regressions algorithm to use based on their performance on standardized data. Instead, a more prudent approach is to consider a variety of algorithms.
Overall, we find that most of the benched algorithms do rather poorly in the benchmark and suggest possible ways to proceed with developing algorithms that will be better at identifying ground truth expressions for cosmological datasets.
As part of this publication we introduce our benchmark algorithm cp3-bench which we make publicly available at https://github.com/CP3-Origins/cp3-bench. The philosophy behind cp3-bench is that is should be as user-friendly as possible, available in a ready-to-use format, and allow for easy additions of new algorithms and datasets.
@@ 1 Introduction
A cores aspect of physics is that we are able to use symbolic expressions to describe how the world around us works. It is therefore a significant challenge when we meet a physical system, observable, or an experimental dataset that we cannot describe through mathematical formulas. This problem arises in various places in astrophysics and cosmology where, for instance, the precise halo mass function is unknown [halo_profile], stellar density profiles are estimated based on various approximations but the true underlying relation is yet to be discovered [stellar_profile], a derived expression for the non-linear matter power spectrum does not yet exist [nonlinear_power, nonlinear_power2], and intricacies of general relativity leave useful expressions for observables such as the mean redshift drift [zdrift1, zdrift2] still waiting to be found (the concept of mean redshift drift will be introduced in section 2). When a ground truth expression for an quantity is desired but seems unfeasible to obtain through analytical considerations, one option is to try to use symbolic regression which is a regression method based on machine learning (ML). Symbolic regression is, at least in principle, a powerful tool for identifying unknown relations between physical quantities such as data from astrophysical observations or laboratory experiments. When utilizing symbolic regression algorithms in e.g. cosmology, the approach has generally been to use one or very few specific algorithms. For instance, the work in [zdrift1, zdrift2] utilized the symbolic regression algorithm AI Feynman [Feynman1, Feynman2] because of its user-friendliness and because it was developed by physicists with discovering physics formulae in mind, showing an improvement compared to earlier state-of-the-art algorithms. The work presented in [nonlinear_power, nonlinear_power2] instead utilized OPERON [operon] which was chosen due to its speed, memory efficiency and because it is based on genetic algorithms which tend to do well in symbolic regression benchmark studies and competitions (e.g. [benchmark, competition]).
@@
In general, when we choose an algorithm for symbolic regression, we naturally make a decision we expect represents the state-of-the-art for the task. However, symbolic regression algorithms are usually evaluated and compared by studying their efficiency at identifying symbolic expressions corresponding to certain ensembles of datasets such as the Penn Machine Learning Benchmarks [PMLB1, PMLB2], collections of mathematical expressions (see e.g. appendix A of [review]), simple physical laws as in EmpiricalBench of [cranmer2023interpretable], and some even use a dataset based on equations from the Feynman Lectures on Physics [FeynmanLectures], the so-called Feynman Symbolic Regression Database [^1] https://space.mit.edu/home/tegmark/aifeynman.html. Although these ensembles often do include physics datasets (not the least the Feynman Symbolic Regression Database), this does not necessarily mean that the resulting benchmark is particularly enlightening when it comes to the applicability of symbolic regression algorithms to real cosmology/astrophysics symbolic regression tasks. One hindrance is that real world cosmology problems are typically based on very complex physics which means that the corresponding datasets can be difficult to obtain, understand and can end up being complicated from a physical point of view, e.g. including several dependent variables. This type of dataset simply does not seem to be represented significantly in common machine learning/symbolic regression databases.
It is not clear that symbolic regression algorithms optimized with typical machine learning database problems are automatically also optimized for real world cosmology problems. We therefore here present a benchmark of symbolic regression algorithms based on specific cosmological datasets. While typical benchmark tasks can consist of hundreds of datasets, we here restrict the benchmark to only 19 datasets which are, however, genuine (synthetic) cosmological datasets related to and representing genuine symbolic regression tasks in cosmology.
To do the benchmark we have developed a new benchmark tool which we call cp3-bench. The version released together with this paper supports the ten different symbolic regression algorithms which are benchmarked in this paper. We provide cp3-bench in a ready-to-use format for Linux machines where installation and usage is straightforward.
@@
In Section 2 below, we introduce the datasets we used for our benchmark, including an introduction to the cosmological settings they represent. In Section 3 we present the benchmark setup, our benchmark tool cp3-bench, example wrapper Things-to-bench and our results. Section LABEL:sec:Summary provides a summary and discussion of future possible developments.
@@ 2 Cosmological datasets
This section serves to describe the data used in our benchmark. In the first subsection below, we introduce the Hubble relation of standard cosmology. The subsequent subsection is dedicated to the observable known as redshift drift and its Friedmann-Lemaitre-Robertson-Walker (FLRW) limit. In the final two subsections we introduce a specific family of toy cosmological models called two-region models and discuss how to compute the redshift drift as well as a quantity known as cosmic backreaction in these models. The datasets are summarized in Section 2.4 which includes a list of the datasets in Table 1.
@@ 2.1 Hubble data
The Hubble parameter describes the expansion history of the Universe and has great interest in cosmology. Earlier work on symbolic regression such as [physics_motivation, Hubble] has thus fittingly included showcasing the use of new symbolic regression algorithms for astrophysics by applying them to Hubble data either directly, or indirectly through supernova data. We will also use Hubble datasets here.
@@
Using this relation, we have generated three different Hubble datasets. The first two are similar to those used in [physics_motivation, Hubble] since we only consider a single cosmological model, namely the (flat) $\Lambda$CDM model with $H_{0}=70$km/s/Mpc and $\Omega_{m,0}=0.3$. One dataset will consist of $(H(z),z)$ without noise while the second dataset will contain 10% Gaussian noise on $H(z)$ to see how big an impact noise has on the performance of the algorithms. For each dataset we compute 1000 data points distributed equidistantly in the redshift interval $z\in[0.1,2]$, roughly representing the redshift interval of real datasets that directly measure the Hubble parameter (e.g. cosmic chronometers [Jimenez_2002, Renzini_2006, Moresco_2018, Moresco_2020]).
We will additionally consider a dataset that corresponds to several $\Lambda$CDM models where we let $H_{0}$ and $\Omega_{m,0}$ vary in the intervals $H_{0}\in[20,100]$ in units of km/s/Mpc and $\Omega_{m,0}\in[0.1,0.5]$. Again, we use equidistant points, now with 50 points for each feature. This dataset, as well as those discussed in the following subsections, are motivated by the ambition to decipher which symbolic regression methods are most suitable for discovering hitherto unknown observational relations. The more general an observational relation is, the more interesting it is. It is therefore desirable to reveal observational relations that are valid for not just a single $\Lambda$CDM model but rather for, say, all FLRW models or $\Lambda$CDM models. We therefore wish to test how well symbolic regression methods can describe data representing one or more families of cosmological models such as sub-families of FLRW models rather than just a single FLRW model. Since this type of dataset will be inherently more complicated than datasets based on only a single FLRW model [^2] Note for instance that it was in [zdrift1, zdrift2] found that the symbolic regression algorithm AI Feynman [Feynman1, Feynman2] was much better at finding analytical expressions for the redshift drift of a single cosmological model than for several models at once., we will not add noise to these datasets.
@@ 2.2 Redshift drift in standard cosmology
where $H(z)$ is the Hubble parameter, $H_{0}$ the Hubble constant and $\delta t_{0}$ is the time interval of the observation such that an object observed to have redshift $z$ at time $t_{0}$ will be observed to have the redshift $z+\delta z$ at time $t_{0}+\delta t_{0}$.
The expression for the redshift written above is fairly simple. Indeed, it represents only a small step in advancement from the direct Hubble data we considered in the previous section. Nonetheless, written as above, it represents an interesting symbolic regression problem because the right hand side of the equation contains a feature $H(z)$ which depends on the other feature ($z$) in a non-trivial way. In addition, $H(z)$ depends on the cosmological model parameters such as $\Omega_{m,0}$ which is not included as a parameter in the dataset since $z$ and $H(z)$ are enough to obtain the expression for the redshift drift shown above. Earlier studies into machine and deep learning algorithms have included datasets where two or more variables were dependent, but these dependencies are typically linear or otherwise simple (see e.g. [BIMT] for an example). However, to the authors’ knowledge, there have not been conducted systematic studies of the impact of the interdependency of features nor studies based on datasets where features have more complicated dependencies such as in the case of the Hubble parameter and its dependence on the redshift.
In cosmology, observable relations are often more convenient to consider in terms of several dependent features. Besides the redshift drift, this is for instance the case for redshift-distance relations which are usually written as an integral over the Hubble parameter in terms of the redshift multiplied by factors containing the redshift itself as well. In our benchmark dataset we will therefore include the redshift drift as written above, i.e. with the target being $\delta z/\delta t_{0}$ and the features being $z$ and $H(z)$. However, we will consider only flat, single-component FLRW models when generating the datasets. In this case, a simple rewriting gives that the redshift drift can be written as
@@
Using the formalism discussed above, we have generated two redshift drift datasets, $(\delta z/\delta t_{0},z,H(z))$ and $(\delta z/\delta t_{0},z,\omega)$. The two models contain the same number of features, has the same target and contains the same number of data points generated using 50 equidistant points in each of the two intervals $z\in[0.1,1]$ and $\omega\in[-0.99,1/3]$. When choosing the redshift interval, we deliberately avoid the very lowest redshifts to avoid a target very close to zero. We choose the upper limit $z=1$ since this is the upper limit expected for main redshift drift measurements from SKA [SKA]. For $\omega$ we choose the upper limit of $1/3$ which corresponds to the equation-of-state parameter for radiation while we choose a lower limit close to that of a cosmological constant, although neglecting the value $-1$ to keep the overall form of the Hubble parameter the same for all models [^4] Remember that for flat, single component universes, $H(z)=\frac{2}{3(1+\omega)t}$ except for the case $\omega=-1$ where the Hubble parameter is constant..
@@ 2.3 Cosmic backreaction and two-region models
The expression for the redshift drift given in the previous section is valid only for FLRW spacetimes, i.e. in spacetimes which have no structures on any scales. Standard cosmology is based on the assumption that the large-scale dynamics of the Universe can be approximated well by the FLRW models but it has also been discussed whether the structures on smaller scales can affect the large-scale/averaged dynamics. This possibility is known as cosmic backreaction [bc_review1, bc_review2, bc_review3, bc_review4] and it has even been suggested that the seemingly accelerated expansion of the Universe is an artifact due to this phenomenon (the co-called backreaction conjecture – see e.g. [conjecture1, conjecture2, conjecture3, conjecture4]). Backreaction and the backreaction conjecture are most commonly considered through the Buchert equations presented in [dust, perfect_fluid, general] where dynamical equations describing the average (large-scale) evolution of the Universe are obtained by averaging Raychaudhuri’s equation and the Hamiltonian constraint using the averaging definition
@@
This equation ensures that the two Buchert equations are consistent with each other. However, the two Buchert equations and the integrability condition do not form a closed set so cannot be used to predict backreaction, not even qualitatively. Therefore, it is unknown how backreaction is realistically parameterized in terms of the scale factor (or equivalently, the mean redshift $\expectationvalue{z}+1\approx 1/a_{D}$ [light1, light2], where triangular brackets indicate taking the mean over many lines of sight to remove statistical fluctuations). As discussed in e.g. [zdrift1, zdrift2], this means that backreaction cannot be sensibly constrained with current cosmological data. This motivated an attempt (presented in [zdrift1, zdrift2]) to obtain parameterizations for backreaction using toy cosmological models where backreaction can be computed, combined with symbolic regression to identify symbolic expression for backreaction. Building on that study, our benchmark dataset includes several datasets based on the so-called two-region models introduced in [2region1, 2region2].
@@
The parameters $f_{u}$ and $f_{o}$ denote the present-time relative volume fractions of the two regions in the ensemble and are related by $f_{u}=1-f_{o}$. As in [2region1, 2region2], present time is defined to correspond to $\phi_{0}=3/2\pi$. As noted in e.g. [zdrift2], once $f_{u}$ or $f_{o}$ has been fixed, there is still one parameter which needs to be set in order to specify a two-region model uniquely. This is seen by noting that $H_{D}=H_{u}(1-v+vh)$, where $v$ is the time dependent volume fraction of the overdense region and $h:=H_{o}/H_{u}$. Since $H_{u}=1/t$ we therefore have $t_{0}=(1-v_{0}+v_{0}h_{0})/H_{D_{0}}$ and thus need to fix either $t_{0}$ or $H_{D_{0}}$. We set $H_{D_{0}}=70$km/s/Mpc.
Since both $Q_{D}$ and $R_{D}$ have unknown parameterizations in terms of $a_{D}$, we have used two types of two-region datasets for the benchmark, namely with $Q_{D}$ and $R_{D}$ as target, respectively. We first consider only a single two-region model specified by $f_{o}=0.2$, yielding the datasets $(Q_{D},\expectationvalue{z})$ and $(R,\expectationvalue{z})$. We then vary $f_{o}$ as well as $z$, leading to the datasets $(Q,\expectationvalue{z},f_{o})$ and $(R_{D},\expectationvalue{z},f_{o})$. Since $f_{o}$ cannot easily be related to FLRW-quantities and hence quantities we often seek to constrain through observations, we also consider the datasets $(Q_{D},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0},H_{D_{0}})$ and $(R_{D},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0},H_{D_{0}})$, where we have defined
@@
Based on the findings in [zdrift1, zdrift2], the data points were generated using equidistant points in the intervals $\expectationvalue{z}\in[0.1,1]$ and $f_{o}\in[0.1,0.25]$.
Since the ground truths of the symbolic expressions for $Q_{D}$ and $R_{D}$ in terms of the features in these datasets are unknown, we asses the success of the symbolic regression algorithms by 1) evaluating their precision on the training model subspace using the mean squared error (MSE), and 2) evaluating their precision on data generated outside the model subspace used for training. Specifically, if a symbolic expression reproduced the data well within the training region, we also considered its performance on a test dataset corresponding to models in the larger interval $z\in[0,5]$ and $f_{o}\in[0.05,0.3]$. The reasoning here is that if a symbolic expression represents the ground truth, it must also be able to reproduce data outside the region it was trained on.
As detailed in [2region1, 2region2], the averaged deceleration parameter $q_{D}$ can be written as

### v2

# cp3-bench: A tool for benchmarking symbolic regression algorithms demonstrated with cosmology

@@ Abstract
We introduce cp3-bench, a tool for comparing/benching symbolic regression algorithms, which we make publicly available at https://github.com/CP3-Origins/cp3-bench. In its current format, cp3-bench includes 12 different symbolic regression algorithms which can be automatically installed as part of cp3-bench. The philosophy behind cp3-bench is that it should be as user-friendly as possible, available in a ready-to-use format, and allow for easy additions of new algorithms and datasets. Our hope is that users of symbolic regression algorithms can use cp3-bench to easily install and compare/bench an array of symbolic regression algorithms to better decide which algorithms to use for their specific tasks at hand.
To introduce and motivate the use of cp3-bench we present a small benchmark of 12 symbolic regression algorithms applied to 28 datasets representing six different cosmological and astroparticle physics setups. Overall, we find that most of the benched algorithms do rather poorly in the benchmark and suggest possible ways to proceed with developing algorithms that will be better at identifying ground truth expressions for cosmological and astroparticle physics datasets. Our demonstration benchmark specifically studies the significance of dimensionality of the feature space and precision of datasets. We find both to be highly important for symbolic regression tasks to be successful. On the other hand, we find no indication that inter-dependence of features in datasets is particularly important, meaning that it is not in general a hindrance for symbolic regression algorithms if datasets e.g. contain both $z$ and $H(z)$ as features. Lastly, we note that we find no indication that performance of algorithms on standardized datasets are good indicators of performance on particular cosmological and astrophysical datasets. This suggests that it is not necessarily prudent to choose symbolic regression algorithms based on their performance on standardized data. Instead, a more robust approach is to consider a variety of algorithms, chosen based on the particular task at hand that one wishes to apply symbolic regression to.
@@ 1 Introduction
A core aspect of physics is that we are able to use symbolic expressions to describe how the world around us works. It is therefore a significant challenge when we meet a physical system, observable, or an experimental dataset that we cannot describe through mathematical formulas. This problem arises in various places in astrophysics and cosmology where, for instance, the precise halo mass function is unknown [halo_profile], stellar density profiles are estimated based on various approximations but the true underlying relation is yet to be discovered [stellar_profile], a derived expression for the non-linear matter power spectrum does not yet exist [nonlinear_power, nonlinear_power2], and intricacies of general relativity leave useful expressions for observables such as the mean redshift drift [zdrift1, zdrift2] still waiting to be found (the concept of mean redshift drift will be introduced in section 2). When a ground truth expression for a quantity is desired but seems unfeasible to obtain through analytical considerations, one option is to try to use symbolic regression which is a regression method based on machine learning (ML). Symbolic regression (SR) is, at least in principle, a powerful tool for identifying unknown relations between physical quantities such as data from astrophysical observations or laboratory experiments.
@@
When utilizing SR algorithms in e.g. cosmology, the approach has generally been to use one or very few specific algorithms. For instance, the work in [zdrift1, zdrift2] utilized the symbolic regression algorithm AI Feynman [Feynman1, Feynman2] because of its user-friendliness and because it was developed by physicists with discovering physics formulae in mind, showing an improvement compared to earlier state-of-the-art algorithms when applied to a specific suite of datasets. The work presented in [nonlinear_power, nonlinear_power2] instead utilized Operon [operon] which was chosen due to its speed, memory efficiency and because it is based on genetic algorithms which tend to do well in SR benchmark studies and competitions (e.g. [benchmark, competition]). In general, when we choose an algorithm for SR, we naturally make a decision we expect represents the state-of-the-art for the task. However, SR algorithms are usually evaluated and compared by studying their efficiency at identifying symbolic expressions corresponding to certain ensembles of datasets such as the Penn Machine Learning Benchmarks [PMLB1, PMLB2], collections of mathematical expressions (see e.g. appendix A of [review]), simple physical laws as in EmpiricalBench of [cranmer2023interpretable], and some even use a dataset based on equations from the Feynman Lectures on Physics [FeynmanLectures], the so-called Feynman Symbolic Regression Database [^1] https://space.mit.edu/home/tegmark/aifeynman.html. The approach of evaluating SR algorithms based on standard datasets seems prudent at least for academic evaluations/comparisons in relation to developing SR algorithms. However, although these ensembles of generic datasets often do include physics datasets (not the least the Feynman Symbolic Regression Database), this does not necessarily mean that the resulting benchmark/comparison is representative of how well the individual algorithms will perform in practice on a specific real cosmology or other (astro-)physics dataset. One hindrance is that real world cosmology problems are typically based on very complex physics which means that the corresponding datasets can be difficult to obtain, understand and can end up being complicated from a physical point of view, e.g. including several dependent variables. This type of dataset simply does not seem to be represented significantly in common machine learning/symbolic regression databases. In addition, it is not clear that SR algorithms optimized with typical machine learning database problems are automatically also optimized for real world physics problems. This means that when one has a genuine SR task, one cannot expect to identify the most optimal algorithms and hyperparameters simply by looking at results from earlier benchmarks or competitions.
To promote an easier and more diverse use of SR algorithms in especially cosmology and astro(particle) physics and to encourage researchers who wish to use but not necessarily develop SR algorithms, we here introduce cp3-bench which we release together with this paper. We have developed cp3-bench as a Python-based tool for comparing/benching SR algorithms. The current version supports 12 different SR algorithms. We provide cp3-bench in a ready-to-use format for Ubuntu machines with Python 3.10 or newer, where installation and usage is straightforward and where the user can choose to have (some or all of) the 12 SR algorithms currently supported automatically installed when installing cp3-bench without extra effort from the user. A true out-of-the-box experience can also be obtained with our Docker image which ensures that the package is ready-to-use on (nearly) any type of machine with Docker and that all dependencies are set correctly.
To demonstrate the use of cp3-bench and to highlight strengths and weaknesses of the 12 SR algorithms as well as obstacles related to different types of datasets, we here present a small benchmark of the 12 SR algorithms with 28 datasets. We stress that the benchmark we present should not itself be used as a guide for deciding which SR algorithm(s) to use for a specific SR task at hand. Instead, our benchmark is meant as a demonstration, and we encourage all to make their own benchmark/comparison of existing SR algorithms to supplement existing “academic” benchmarks such as [benchmark, competition] when choosing which SR algorithms to use for practical tasks at hand. We encourage contributions to the benchmark e.g. by adding new algorithms to the suite.
@@
In Section 2 below, we introduce the datasets we used in our benchmark, including an introduction to the physical settings they represent. In Section 3 we present the benchmark setup, our benchmark tool cp3-bench and an example wrapper Things-to-bench. Section 3.2 presents our benchmark results. Section LABEL:sec:Summary provides a summary and discussion of future possible developments.
@@ 2 Cosmological datasets
This section serves to describe the data used in our benchmark. There can be several different aims when using SR. A goal can for instance be to make a good fit of the nonlinear matter spectrum for certain $\Lambda$CDM model parameters and use the fit as a type of emulator. Another goal can be to identify new laws of nature or hitherto unknown observational relations, i.e. discovering “ground truth” expressions. We will focus our demonstration benchmark on the latter and our choice of benchmark data and evaluation criteria of SR results are based on having this goal in mind. Note that most of the synthetic datasets we choose do not themselves represent actual SR problems, i.e. for most of the datasets we consider, the ground truth is known. This makes it easier to evaluate the performance of SR algorithms in terms of their ability to specifically “guess” ground truths. The main differences here compared to standard datasets are thus that a) the data is based on real cosmology and astro(particle) physics models, making it more closely related to datasets one could imagine considered in real astropysics applications, and b) we specifically choose datasets and variations of these that highlight different qualities of datasets that may be important when seeking to use SR to identify ground truth expressions. In addition, two groups of our datasets do indeed represent datasets where the ground truth is unknown.
Below, we introduce six overall groups of datasets. Each group of datasets represents a specific physical scenario and within each group there can be several versions of datasets representing the scenario. In the first subsection below, we introduce our first group of datasets which is based on the Hubble relation of standard cosmology. The following subsection is dedicated to the observable known as redshift drift and its Friedmann-Lemaitre-Robertson-Walker (FLRW) limit. In the subsequent two subsections we introduce a specific family of toy cosmological models called two-region models and discuss how to compute the redshift drift as well as a quantity known as cosmic backreaction in these models. We then introduce datasets based on dark matter galaxy halo profiles and lastly a group of datasets based on simplified mock gravitational wave data. The datasets are summarized in Section 2.6 which includes a list of the datasets in Table 1.
Since most of the 28 datasets are cosmology-based, we will for simplicity refer to this group of datasets as the “cosmological” datasets to distinguish them from a group of test datasets introduced later.
Figure 1: The left figure shows $C1a$ ($H(z)$ with no error for a single $\Lambda$CDM model) and $C1b$ ($H(z)$ with 10 % Gaussian error for a single $\Lambda$CDM model). The figure to the right shows the datasets $C1c$ and $C1d$, i.e. $H(z)$ for $\Lambda$CDM models with varying $\Omega_{m}$ and both $\Omega_{m}$ and $H_{0}$, respectively. For $C1c$ and $C1d$, one can see the dependence of $\Omega_{m}$ (and $H_{0}$ in the latter) as vertical lines in the dataset at constant $z$.
@@ 2.1 Hubble data
The Hubble parameter describes the expansion history of the Universe and has great interest in cosmology. Earlier work on symbolic regression such as [physics_motivation, Hubble, language] has thus fittingly included showcasing the use of new symbolic regression algorithms for astrophysics by applying them to Hubble data either directly, or indirectly through supernova data. We will also use Hubble datasets here.
@@
Using this relation, we have generated three different Hubble datasets. The first two are similar to those used in [physics_motivation, Hubble] since we only consider a single cosmological model, namely the (flat) $\Lambda$CDM model with $H_{0}=70$km/s/Mpc and $\Omega_{m,0}=0.3$. One dataset will consist of $(H(z),z)$ without noise while the second dataset will contain 10% Gaussian noise on $H(z)$. By including datasets that are identical up to noise we can asses the impact that noise has on the performance of the SR algorithms. For each dataset we compute 1000 data points distributed equidistantly in the redshift interval $z\in[0.1,2]$, roughly representing the redshift interval of real datasets that directly measure the Hubble parameter (e.g. measurements based on cosmic chronometers [Jimenez_2002, Renzini_2006, Moresco_2018, Moresco_2020] or on baryon acoustic oscillations, BAO [BAO_1, BAO_2, BAO_3, BAO_4]). We will refer to these two datasets as $C1a$ and $C1b$. They are plotted together in Figure 1.
We will additionally consider a dataset that corresponds to several $\Lambda$CDM models where we let $H_{0}$ and $\Omega_{m,0}$ vary in the intervals $H_{0}\in[20,100]$ in units of km/s/Mpc and $\Omega_{m,0}\in[0.1,0.5]$. Again, we use equidistant points, now with 50 points for each feature. This dataset, as well as those discussed in the following subsections, are motivated by the ambition to decipher which symbolic regression methods are most suitable for discovering hitherto unknown observational relations. The more general an observational relation is, the more interesting it is. It is therefore desirable to reveal observational relations that are valid for not just a single $\Lambda$CDM model but rather for, say, all FLRW models or $\Lambda$CDM models. We therefore wish to test how well symbolic regression methods can describe data representing one or more families of cosmological models such as sub-families of FLRW models rather than just a single FLRW model. Since this type of dataset will be inherently more complicated than datasets based on only a single FLRW model [^2] Note for instance that it was in [zdrift1, zdrift2] found that the symbolic regression algorithm AI Feynman [Feynman1, Feynman2] was much better at finding analytical expressions for the redshift drift of a single cosmological model than for several models at once., we will not add noise to these datasets. We denote these two datasets $C1c$ and $C1d$, respectively, and show them in Figure 1.
Figure 2: The redshift drift datasets $C2a$ and $C2b$ plotted against the different features used in the two datasets.
@@ 2.2 Redshift drift in standard cosmology
where $H(z)$ is the Hubble parameter, $H_{0}$ the Hubble constant, and $\delta t_{0}$ is the time interval of the observation such that an object observed to have redshift $z$ at time $t_{0}$ will be observed to have the redshift $z+\delta z$ at time $t_{0}+\delta t_{0}$.
The expression for the redshift drift written above is fairly simple. Nonetheless, written as above, it represents an interesting symbolic regression problem because the right hand side of the equation contains a feature ($H(z)$) which depends on the other feature ($z$) in a non-trivial way. In addition, $H(z)$ depends on the cosmological model parameters such as $\Omega_{m,0}$ which is not included as a parameter in the dataset since $z$ and $H(z)$ are enough to obtain the expression for the redshift drift shown above. Earlier studies into machine and deep learning algorithms have included datasets where two or more variables were dependent, but these dependencies are typically linear or otherwise simple (see e.g. [BIMT] for an example). However, to the authors’ knowledge, there have not been conducted systematic studies of the impact of the inter-dependence of features nor studies based on datasets where features have more complicated dependencies such as in the case of the Hubble parameter and its dependence on the redshift.
In cosmology, observable relations are often more convenient to consider in terms of several dependent features. Besides the redshift drift, this is for instance the case for redshift-distance relations which are usually written as integrals over the Hubble parameter in terms of the redshift multiplied by factors containing the redshift itself as well. In our benchmark dataset we will therefore include the redshift drift as written above, i.e. with the target being $\delta z/\delta t_{0}$ and the features being $z$ and $H(z)$. However, we will consider only flat, single-component FLRW models when generating the datasets. In this case, a simple rewriting of the above expression yields
@@
Using the formalism discussed above, we have generated two redshift drift datasets, $(\delta z/\delta t_{0},z,H(z,\omega))$ and $(\delta z/\delta t_{0},z,\omega)$. The two models contain the same number of features, has the same target and contains the same number of data points generated using 50 equidistant points in each of the two intervals $z\in[0.1,1]$ and $\omega\in[-0.99,1/3]$, resulting in a total of $50^{2}$ data points per dataset. When choosing the redshift interval, we deliberately avoid the very lowest redshifts to avoid a target very close to zero. This is sensible since some of the SR algorithms we bench (e.g. AI Feynman) remove data with low values of the target which we wish to avoid. We choose the upper limit $z=1$ since this is the upper limit expected for main redshift drift measurements from SKA [SKA]. For $\omega$ we choose the upper limit of $1/3$ which corresponds to the equation-of-state parameter for radiation while we choose a lower limit close to that of a cosmological constant, although neglecting the value $-1$ to keep the overall form of the Hubble parameter the same for all models [^4] Remember that for flat, single component universes, $H(z)=\frac{2}{3(1+\omega)t}$ except for the case $\omega=-1$ where the Hubble parameter is constant.. We denote the redshift drift datasets $C1a$ and $C1b$ and plot the redshift drift against $z,H(z)$ and $\omega$ in Figure 2.
Figure 3: Left: $Q_{D}$ and $R_{D}$ plotted against $z$, corresponding to the datasets $C3a-C3d$. Middle: $Q_{D}$ plotted against different features (dataset $C3e$). Right: $R_{D}$ plotted against different features (dataset $C3f$).
@@ 2.3 Cosmic backreaction and two-region models
The expression for the redshift drift given in the previous section is valid only for FLRW spacetimes, i.e. in spacetimes which have no structures on any scales. Standard cosmology is based on the assumption that the large-scale dynamics of the Universe can be approximated well by the FLRW models but it has also been discussed whether the structures on smaller scales can affect the large-scale/averaged dynamics. This possibility is known as cosmic backreaction [bc_review1, bc_review2, bc_review3, bc_review4] and it has even been suggested that the seemingly accelerated expansion of the Universe is an artifact due to this phenomenon (the co-called backreaction conjecture – see e.g. [conjecture1, conjecture3, conjecture4]). Backreaction and the backreaction conjecture are most commonly considered through the Buchert equations presented in [dust, perfect_fluid, general] where dynamical equations describing the average (large-scale) evolution of the Universe are obtained by averaging Raychaudhuri’s equation and the Hamiltonian constraint using the averaging definition
@@
This equation ensures that the two Buchert equations are consistent with each other. However, the two Buchert equations and the integrability condition do not form a closed set so cannot be used to predict backreaction, not even qualitatively. Therefore, it is unknown how backreaction is realistically parameterized in terms of the scale factor (or equivalently, the mean redshift $\expectationvalue{z}+1\approx 1/a_{D}$ [light1, light2], where triangular brackets indicate taking the mean over many lines of sight to remove statistical fluctuations). As discussed in e.g. [zdrift1, zdrift2], this means that backreaction cannot be sensibly constrained with current cosmological data. This motivated an attempt (presented in [zdrift1, zdrift2]) to obtain parameterizations for backreaction using toy cosmological models where backreaction can be computed, combined with symbolic regression to identify symbolic expressions for backreaction. Building on that study, our benchmark dataset includes several datasets based on the so-called two-region models introduced in [2region1, 2region2].
@@
The parameters $f_{u}$ and $f_{o}$ denote the present-time relative volume fractions of the two regions in the ensemble and are related by $f_{u}=1-f_{o}$. As in [2region1, 2region2], present time is defined to correspond to $\phi_{0}=3/2\pi$. As noted in e.g. [zdrift2], once $f_{u}$ or $f_{o}$ has been fixed, there is still one parameter which needs to be set in order to specify a two-region model uniquely. This is seen by noting that $H_{D}=H_{u}(1-v+vh)$, where $v$ is the time-dependent volume fraction of the overdense region and $h:=H_{o}/H_{u}$. Since $H_{u}=1/t$ we therefore have $t_{0}=(1-v_{0}+v_{0}h_{0})/H_{D_{0}}$ and thus need to fix either $t_{0}$ or $H_{D_{0}}$. We set $H_{D_{0}}=70$km/s/Mpc.
Since both $Q_{D}$ and $R_{D}$ have unknown parameterizations in terms of $a_{D}$, we have used two types of two-region datasets for the benchmark, namely with $Q_{D}$ and $R_{D}$ as target, respectively. We first consider only a single two-region model specified by $f_{o}=0.2$, yielding the datasets $(Q_{D},\expectationvalue{z})$ and $(R_{D},\expectationvalue{z})$. We denote these two datasets $C3a$ and $C3b$, respectively. We then vary $f_{o}$ as well as $z$, leading to the datasets $C3c=(Q_{D},\expectationvalue{z},f_{o})$ and $C3d=(R_{D},\expectationvalue{z},f_{o})$. Since $f_{o}$ cannot easily be related to FLRW-quantities and hence quantities we often seek to constrain through observations, we also consider the datasets $C3e=(Q_{D},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0},H_{D_{0}})$ and $C3d=(R_{D},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0},H_{D_{0}})$, where we have defined
@@
Based on the findings in [zdrift1, zdrift2], the data points were generated using equidistant points in the intervals $\expectationvalue{z}\in[0.1,1]$ and $f_{o}\in[0.1,0.25]$ with $50$ points in both dimensions (giving $50^{2}$ points in total per dataset). We note that $H_{D_{0}}$ is included as a feature in these datasets even though it takes the same value in all data points. Ideally, SR algorithms should not include this feature in their results. The datasets are shown in Figure 3.
Since the ground truths of the symbolic expressions for $Q_{D}$ and $R_{D}$ in terms of the features in these datasets are unknown, we asses the success of the symbolic regression algorithms by 1) evaluating their precision on the training subspace using the mean squared error (MSE), and 2) evaluating their precision on data generated outside the parameter subspace used for training. Specifically, if a symbolic expression reproduces the data well within the training region, we also consider its performance on a test dataset corresponding to models in the larger interval $z\in[0,5]$ and $f_{o}\in[0.05,0.3]$. The reasoning here is that if a symbolic expression represents the ground truth, it must also be able to reproduce data outside the region it was trained on. Evaluating the performance of SR algorithms based on the ability of the resulting symbolic expressions to extrapolate well outside the training region is only sensible because we are specifically interested in the ability of SR algorithms to identify ground truth expression which must naturally extrapolate well outside the training region. If we were merely interested in good fits/emulators within a specific region, a more prudent approach would be to make training data within the entire desired region and only evaluate performance based on this region.
Figure 4: Left: $C3g$ shown as $Q_{D}$ plotted against the different features of the dataset. Right: $C3h$ shown as $Q_{D}$ plotted against the different features of the dataset.
The datasets $C3e$ and $C3f$ are plotted in Figure 3. We note that the plots of $C3e$ and $C3f$ show $Q_{D}$ and $R_{D}$ to be close to linear in $\Omega_{m,0},\Omega_{R,0}$ and $\Omega_{Q,0}$ which would indicate it reasonable to expect the SR algorithms to identify good fits within the training region (which might or might not represent ground truths).
As detailed in [2region1, 2region2], the averaged deceleration parameter of two-region models, $q_{D}$, can be written as